• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jack (Techs)

    Log entry time 19:00:42 on October 06, 2010

    Entry number 333831

    keyword=Magnet testing status

    The LQ1 has not yet been tested. As noted in Howard's previous entry, he is in the midst of changing the SCRs of LQ2. LQ2 looked good till it would trip off from over temperature. So, I expect the SCRs will be the only issue with the LQ2. The LQ3 looked good all the way to 1600A. So, I think the LQ3 can be considered ready. We have left it off since cryogenics are still varying a bit and the LQ3 needs a lot of leadflow to run. The Left Dipole was run up to 3.5GeV and sat there for quite a bit while we looked at it. The magnet and power supply appear to be usable. Some of the scope pictures off the power supply look a little suspicious, but that needs to be looked at later and would only be routine maintenance concerns. However, we found the NMR would not lock at 3.5GeV. Investigating further, the indications are that the multiplexer inside the dipole only functions on the A probe. The A probe covers the lowest momentum settings of the NMR.